OpenStudy (anonymous):

An object is launched at 19.6 meters per second (m/s) from a 58.8-meter tall platform. The equation for the object's height s at time t seconds after launch is s(t) = –4.9t2 + 19.6t + 58.8, where s is in meters. When does the object strike the ground?

OpenStudy (anonymous):

0 = (t -6)(t + 2) => t=6 and t=-2 Correct solution is t=6 because negative time is not possible

OpenStudy (anonymous):

Erm.. I got 1.29,-9.29

OpenStudy (anonymous):

you can rewrite the equation as 0 =t2 ± 4t ± 12

OpenStudy (anonymous):

You're right mkuma.. I just scratched it out :/

OpenStudy (anonymous):

t^2 - 4t -12 =0 (t-4)*(t+3) =0